### v1

# CausalRAG: Integrating Causal Graphs into Retrieval-Augmented Generation

Abstract
Large language models (LLMs) have revolutionized natural language processing (NLP), particularly through Retrieval-Augmented Generation (RAG), which enhances LLM capabilities by integrating external knowledge. However, traditional RAG systems face critical limitations, including disrupted contextual integrity due to text chunking, and over-reliance on semantic similarity for retrieval. To address these issues, we propose CausalRAG, a novel framework that incorporates causal graphs into the retrieval process. By constructing and tracing causal relationships, CausalRAG preserves contextual continuity and improves retrieval precision, leading to more accurate and interpretable responses. We evaluate CausalRAG against regular RAG and graph-based RAG approaches, demonstrating its superiority across several metrics. Our findings suggest that grounding retrieval in causal reasoning provides a promising approach to knowledge-intensive tasks.
CausalRAG: Integrating Causal Graphs into Retrieval-Augmented Generation
Nengbo Wang1,2, Xiaotian Han1, Jagdip Singh2, Jing Ma1, Vipin Chaudhary1 1Department of Computer and Data Sciences, Case Western Reserve University 2Department of Design and Innovation, Case Western Reserve University nengbo.wang@case.edu
1 Introduction
The rapid advancements in large language models (LLMs) have revolutionized the field of natural language processing (NLP), enabling a wide range of applications Anthropic (2024); Google (2024); OpenAI (2024). However, their reliance on pre-trained knowledge limits their ability to integrate and reason over dynamically updated external information, particularly in knowledge intensive domains such as academic research. Retrieval-Augmented Generation (RAG) has emerged as a promising framework to address this limitation Lewis et al. (2021), combining retrieval mechanisms with generative capabilities to enhance contextual understanding and response quality.
Recent research has focused on improving RAG along two primary directions: 1) enhancing retrieval efficiency and integration mechanisms by designing more adaptive and dynamic retrieval frameworks Gan et al. (2024); Ravuru et al. (2024); Zhang et al. (2024a); 2) improving the representation of external knowledge to facilitate retrieval and reasoning, with graph-based RAGs being a dominant approach Edge et al. (2024); Guo et al. (2024); Potts (2024). Despite these advancements, existing RAG architectures still face critical limitations that impact retrieval quality and response accuracy, primarily due to three key issues: 1) disruption of contextual integrity caused by the text chunking design; 2) reliance on semantic similarity rather than causal relevance for retrieval; and 3) a lack of accuracy in selecting truly relevant documents.
Through a combination of theoretical analysis and empirical evaluation, we rethink the limitations of current RAG systems by introducing a novel perspective based on context recall and precision metrics. Our findings reveal that both regular and graph-based RAGs struggle not only to retrieve truly grounded context but also to accurately discern the relationship between retrieved content and the user query. We identify this fundamental issue as one primary reason why LLMs in RAG frameworks often generate seemingly relevant yet shallow responses that lack essential details.
To address these gaps, we introduce CausalRAG, a novel RAG framework that integrates causal graphs to enhance retrieval accuracy and reasoning performance. Unlike regular and graph-based RAGs, CausalRAG constructs a causal graph from uploaded documents, preserving contextual relationships while capturing cause-effect dependencies. By ensuring that retrieved documents are both relevant and causally grounded, CausalRAG enables the generation of more contextually rich and causally detailed responses. This approach not only improves retrieval effectiveness but also mitigates hallucinations and enhances answer faithfulness.
We evaluate CausalRAG on datasets from diverse domains and across varying context lengths, comparing its performance with regular RAG and GraphRAG, a graph-based RAG framework from Microsoft Edge et al. (2024). Our experiments assess performance across three key metrics: answer faithfulness, context recall, and context precision. Results demonstrate that CausalRAG achieves superior performance across different contexts. Additionally, we conduct a case study and a parameter analysis to further examine our framework, analyzing and providing insights that contribute to ongoing research in RAG. The contributions of this work are threefold:
  • We systematically identify the inherent limitations of RAG’s retrieval process through analytical and experimental study. More importantly, we uncover one major reason why LLMs in RAG tend to generate superficial, generalized answers that lack the grounded details expected by users.
  • We propose CausalRAG, a framework that enhances both retrieval and generation quality by incorporating causality into the RAG, effectively addressing these limitations.
  • Our work further mitigates hallucination issues and significantly improves the interpretability of AI systems. We summarize key findings and insights in both retrieval and generation process, contributing to research in RAG.
2 Related Work
2.1 Retrieval-Augmented Generation
RAG enhances LLMs’ ability to handle knowledge-intensive tasks by integrating external knowledge retrieval Lewis et al. (2021). Existing research primarily advances RAG along two key dimensions: 1) improving retrieval efficiency and integration mechanisms; 2) enhancing the representation of external knowledge to facilitate reasoning and interpretability.
Optimizing retrieval flow and interaction. The first stream focuses on improving the interaction flow within the RAG system to enhance output quality. Approaches have introduced pre-retrieval, retrieval, and post-retrieval refinements to mitigate redundancy and computational overhead Wang et al. (2024). Modular RAG architectures further advance this by enabling iterative retrieval-generation cycles, allowing dynamic interactions between retrieval and content creation. For example, CAiRE-COVID Su et al. (2020) demonstrated the effectiveness of iterative retrieval in multi-document summarization, while some work Feng et al. (2023) extended this approach to multi-hop question answering. Recent innovations include METRAG Gan et al. (2024), which integrates LLM supervision to generate utility-driven retrieval processes, and RAFT Zhang et al. (2024a), which trains models to disregard distractor documents while improving citation accuracy through chain-of-thought reasoning.
Structuring external knowledge for efficiency. The second direction explores improved methods for structuring external knowledge to achieve better retrieval efficiency. For example, GraphRAG Edge et al. (2024), proposed by Microsoft, treats external knowledge as interconnected nodes, capturing causal and thematic relationships to enhance retrieval depth and reasoning. LightRAG introduced a dual-level retrieval mechanism for incremental knowledge updates Guo et al. (2024). More recently, Lazy GraphRAG has been developed as a continuous version of GraphRAG, deferring computationally expensive operations until query time and leveraging lightweight indexing techniques for efficiency Potts (2024). Despite these advancements, ensuring the quality and relevance of retrieved documents remains a significant concern in current RAG systems, as it directly impacts the coherence of the final response Gupta et al. (2024).
2.2 Causal Graphs and RAG
The combination of causal graphs and RAG has emerged as a promising approach to enhancing knowledge retrieval and reasoning. As causality provides a structural understanding of dependencies within data, it enables more interpretable and reliable AI outputs Ma (2024). Existing research in this domain primarily advances causal discovery with RAG and LLMs. For instance, some work proposed an LLM-assisted breadth-first search (BFS) method for full causal graph discovery, significantly reducing time complexity Jiralerspong et al. (2024). Additionally, some further introduced a correlation-to-causation inference (Corr2Cause) task to evaluate LLMs’ ability to infer causation from correlation, revealing their limitations in generalization across different datasets (Jin et al., 2024).
Despite these advancements, most studies focus on utilizing RAG or LLMs for causal discovery or causal effect estimation Ma (2024); Kıcıman et al. (2024), whereas the direct integration of causal graphs into RAG architectures remains largely unexplored. Our work aims to be a pioneer in this direction. A few existing studies have touched upon this concept but differ in scope. One approach integrates causal graphs within the LLM architecture itself, structuring the transformer’s internal token processing using causality rather than enhancing RAG retrieval Chen et al. (2024b). Another employs causal graphs in RAG systems but focuses on the pre-retrieval stage and largely reduces the core process into a single embedding model without deeper exploration Samarajeewa et al. (2024). GraphRAG is a well-regarded and influential work in this area, as it introduces a graph-based structure into the RAG system, leveraging graph community detection and summarization techniques for retrieval Edge et al. (2024). While it does not incorporate causality, it has significantly improved RAG performance. Therefore, we adopt GraphRAG as the baseline for our work.
In the following sections, we first analyze the nature of regular RAG and grah-based RAGs from a novel perspective, identifying their inherent limitations. We then introduce a causal graph structure to address these gaps within the RAG system and present our framework—CausalRAG.
Figure 1: Analytical and experimental studies reveal limitations in regular RAG and GraphRAG. (a) identifies three key retrieval and generation issues in regular RAG; (b) evaluates RAG via context precision and recall, showing regular RAG excels in recall but lacks precision. GraphRAG improves precision but trades off some recall.
3 Why Regular RAG Fails in Providing Accurate Responses
In this section, through both analytical and experimental investigations, we identify three fundamental limitations of regular RAG and rethink its design by examining its three core elements—user query, retrieved context, and response—through a novel perspective based on precision and recall.
3.1 Limitations of regular RAG
The first limitation arises from RAG’s common practice of chunking texts into minimal units (as illustrated in Figure 1a). This process disrupts the natural linguistic and logical connections in the original text. These connections are crucial for maintaining contextual integrity, and if they are lost, an alternative mechanism must be implemented to restore them.
The second limitation lies in the semantic search process. RAG typically retrieves the semantically closest documents from a vector database based on query similarity. However, in many cases, critical information necessary for answering a query is not semantically similar but rather causally relevant. A classic example is the relationship between diapers and beer—while they are not semantically related, they may exhibit a causal connection in real worlds. This limitation suggests that RAG’s reliance on semantic similarity may lead to the retrieval of contextually irrelevant but superficially related information.
The third limitation is that even when RAG retrieves a relevant context, this does not necessarily guarantee an accurate response. To formalize this issue, we used two key metrics: context recall and context precision, defined as follows:
Context $\text{Context Recall}=\frac{\sum_{i=1}^{N}\mathbb{I}(C_{i}\in R)}{|R|}$ (1)
Here $R$ is the reference set of all relevant reference documents. $C_{i}$ is the $i^{th}$ retrieved document. $\mathbb{I}(C_{i}\in R)$ is an indicator function that returns 1 if $C_{i}$ belongs to the reference set $R$, otherwise 0. It should always return 1 if no hallucination occurs in the LLM.
Context $\text{Context Precision}=\frac{\sum_{i=1}^{N}\mathbb{I_{Q}}(C_{i}\in R)}{\sum_{i=1}^{N}\mathbb{I}(C_{i}\in R)}$ (2)
Here $\mathbb{I_{Q}}(C_{i}\in R)$ is an indicator function that returns 1 if the context retrieved is causally related to user’s query, otherwise 0.
Recall-Precision Perspective. Context recall measures how much of the correct contextual information can be retrieved from the user’s uploaded documents given a query. While RAG primarily focuses on retrieving related documents, adjusting retrieval parameters to increase the number of retrieved documents can generally lead to higher recall in reality. However, context precision presents a challenge. It measures the proportion of retrieved documents that are actually correct given the user query. As discussed earlier, RAG’s reliance on semantic similarity rather than causal relevance often leads to the retrieval of superficially similar but logically irrelevant content. In summary, while RAG can recall numerous answers from reference materials, the proportion of correct responses remains low, ultimately reducing its precision. This recall-precision perspective provides a new lens to see limitations of regular RAG.
3.2 Rethinking Graph-based RAGs
Applying this perspective, we can better understand why Graph-based RAGs serve as improved variants of RAG. By summarizing and ranking the importance of graph communities before retrieval, they largely enhance the quality of retrieved context, thereby improving context precision. However, it only partially addresses the identified limitations, as its summarization process does not entirely filter out irrelevant information. More critically, its reliance on community-based summarization for retrieval may adversely impact recall. Based on these analytical insights, we further conducted an experimental study to validate our analysis.
Experimental study. As shown in Figure 1(b), we conducted an experimental study to empirically verify our analysis. Specifically, we employed an LLM to evaluate the performance of regular RAG and GraphRAG using the RAG evaluation framework Ragas Es et al. (2023). It takes as input the three key elements—user query, retrieved context, and reference—along with predefined metric definitions, and prompts the LLM to assign numerical ratings. This LLM-based evaluation has been widely used in recent RAG research and has also been adopted in RAG evaluations Samarajeewa et al. (2024); Edge et al. (2024).
The results indicate that both global and local versions of GraphRAG achieve higher context precision compared to regular RAG. However, its recall performance remains unsatisfactory, even slightly lower than that of RAG, due to its graph community-based retrieval process.
By combining our precision-recall-based analytical analysis with empirical findings, we clearly outline the inherent limitations of both RAG and its graph-based extensions. In the next section, we introduce our proposed framework, CausalRAG, designed to address these issues.
Figure 2: Overview of CausalRAG’s architecture. Documents are indexed as graphs, and queries retrieve causally related nodes. A causal summary is generated and combined with the query to ensure grounded responses.
4 Methodology
In this section, we introduce our novel framework—CausalRAG—which integrates RAG with causality to overcome the limitations of existing RAG systems. Overall, CausalRAG utilizes a graph-based approach to represent uploaded documents, enhancing robustness against discontinuities caused by text chunking (as shown in Figure 2). More importantly, by expanding and tracing nodes in the graph through causal relationships, CausalRAG is able to retrieve a more diverse set of contextual information while maintaining causal groundedness. This enables CausalRAG to achieve both high recall and strong precision. We now discuss each step in detail.
4.1 Indexing
At the outset, once the system receives the user’s uploaded documents and query, we first index these inputs into our vector database. For the uploaded documents, we employ a text-based graph construction method to transform the text into a structured graph before storing the nodes and edges in the vector database. Specifically, we leverage an LLM to construct the graph following the approach by LangChain Chase (2022), where the LLM scans the text to identify graph nodes and determine relationships between them. Although LLM-based graph construction has shown strong performance Chen et al. (2024a) and is widely adopted in graph-based RAG research , we further validate our constructed graph using expert knowledge, which we elaborate on in the case study. After constructing and indexing the base graph, we embed the user query into the vector database, preparing it for subsequent search and matching. Notably, this indexing process occurs independently of query time, allowing for efficient retrieval during inference.
4.2 Discovering and Estimating Causal Paths
This is the first step during query time. We begin by matching the user query to nodes in the graph based on their embedding distance. The $k$ nodes with the smallest distances are selected, representing the most relevant information directly related to the query. Notably, $k$ is a tunable parameter—higher values retrieve more relevant information at the cost of increased computational complexity.
After selecting the initial $k$ nodes, we expand the search along the base graph’s edges by a step size of $s$, thereby broadening the retrieved context. This step is crucial as it preserves causal and relational connections within the text, allowing CausalRAG to retrieve more context while maintaining high recall. The parameter $s$ controls the depth of expansion, where higher values lead to more diverse information retrieval.
Once the relevant nodes and edges are collected, we employ an LLM to identify and estimate causal paths within them, constructing a refined causal graph. As discussed, LLMs have demonstrated efficiency in discerning and analyzing causal relationships Zhang et al. (2024b); Zhou et al. (2024). This step ensures that CausalRAG prioritizes causally relevant information, improving precision.
Furthermore, the derived causal graph serves two key purposes: 1) It preserves causally relevant information that traditional retrieval methods struggle to capture. More importantly, by adjusting the parameter $s$, this approach can capture long-range causal relationships within the text, particularly when the text is lengthy; 2) It filters out semantically related but causally irrelevant information. Without this filtering, responses may contain unnecessary or even hallucinated content, compromising answer faithfulness.
4.3 Retrieving Context Causally
After constructing the causal graph, we summarize the retrieved information and generate a causal summary. Notably, the input at this stage consists of information that is not only highly relevant but also causally grounded in the user’s query, ensuring greater validity. This approach contrasts with traditional retrieval methods, which often rely purely on semantic similarity and may retrieve contextually related yet causally irrelevant information.
The causal summary is derived by tracing key causal paths within the graph, prioritizing nodes and relationships that contribute directly to answering the query. This ensures that the retrieved information maintains logical coherence and factual consistency while filtering out spurious or weakly related context. Additionally, by leveraging causal dependencies, our method reduces the risk of retrieving semantically similar but misleading evidence.
Once the causal summary is generated, it is combined with the user query to construct a refined prompt for CausalRAG. This structured final input allows RAG to focus on reasoning through causal relationships rather than merely aggregating loosely related text spans.
5 Experiment
To evaluate the effectiveness of CausalRAG, we conduct a series of experiments comparing it with regular RAG and GraphRAG baselines. Our evaluation spans multiple datasets, retrieval settings, and performance metrics, ensuring a comprehensive analysis of retrieval quality and answer faithfulness. We systematically explore different parameter settings and retrieval strategies to assess their impact on model performance.
5.1 Experimental Setup
Parameters & Baselines. We set the default parameters for CausalRAG as $k=s=3$ and use the same $k$ value for GraphRAG’s community-based retrieval and regular RAG’s document retrieval to ensure fairness. We evaluate four RAG variants: regular RAG Lewis et al. (2021), GraphRAG with local and global search Edge et al. (2024), and our proposed CausalRAG. Local search retrieves from raw document graphs, making it ideal for passage-specific queries, while global search relies on graph community summaries for broader context understanding. We use GPT-4o-mini as the base LLM for all models.
Datasets. We evaluate on publicly available research papers from various domains, ranging from 100 to around 17,000 tokens, reflecting typical user-uploaded documents. An LLM will generate $n$ grounded questions per document, ensuring they are explicitly answerable. We use the Ragas framework Es et al. (2023) and assess models on three key metrics: answer faithfulness, context precision, and context recall. Faithfulness measures factual consistency on a scale from 0 to 1, with higher scores indicating better alignment with reference documents.
5.2 Performance Comparison
To comprehensively evaluate the effectiveness of CausalRAG, we conducted experiments on a diverse dataset comprising abstracts from 100 randomly selected research papers using OpenAlexPriem et al. (2022). All evaluations were performed using the OpenAlex public academic dataset, which spans multiple disciplines, including applied mathematics, art history, library science, and psychology, ensuring a comprehensive assessment across different domains. For each document, we constructed graph and generated $n=5$ questions per paper for assessing retrieval and response quality. As mentioned, we then compared CausalRAG’s performance against three baselines—regular RAG, GraphRAG-Local, and GraphRAG-Global—across three key metrics: answer faithfulness, context recall, and context precision.
Figure 3 presents the comparative results. Overall, CausalRAG consistently outperforms all other models across the three evaluation metrics, demonstrating its ability to generate more factually accurate responses while maintaining both high recall and precision.
Answer Faithfulness. This metric measures the extent to which the generated response aligns with the reference information, ensuring factual correctness. As shown in Figure 3, regular RAG achieves relatively strong performance in this metric, suggesting that despite its reliance on purely semantic retrieval, it still retrieves some relevant context. However, GraphRAG-Local still slightly outperforms regular RAG by leveraging structured graph-based retrieval locally, which supports its grounded answer. Naturally, GraphRAG-Global CausalRAG supported by high-level community summary performs worse in this metric. Lastly, CausalRAG holds its answer faithfulness at a good level by retrieving causally grounded context, reducing hallucinations, and ensuring responses are not only relevant but also justified by retrieved evidence.
Context Recall. Context recall evaluates how much of the correct reference information is retrieved during the RAG process. As expected, regular RAG exhibits high recall, given its broad retrieval approach, which tends to maximize the inclusion of potentially relevant content. GraphRAG-Local and GraphRAG-Global show a slight reduction in recall due to their community-based summarization, which, while improving retrieving process, sacrifices some context diversity. CausalRAG, by contrast, strikes a balance between precision and recall and its recall is still higher than regular RAG. It is ensured by that retrieved information is causally relevant, avoiding excessive retrieval of loosely related but non-causal information.
Context Precision. Precision is where CausalRAG demonstrates the most significant advantage. While regular RAG achieves relatively low precision—often retrieving semantically similar yet contextually irrelevant content—GraphRAG-Local and GraphRAG-Global significantly improve upon this by leveraging graph structures to organize knowledge more effectively. However, both still struggle with retrieving causally relevant content and are lower than CausalRAG on this metric. We can see CausalRAG still maintain a great precision, as it inherently filters retrieved information through causal graph.
Figure 3: Performance comparison of Regular RAG, GraphRAG-Global, GraphRAG-Local, and CausalRAG on the OpenAlex dataset across three key metrics: answer faithfulness, context recall, and context precision.
5.3 Case Study
Figure 4: Case Study comparing retrieving process between different RAG systems.
Figure 5: Performance comparison of four RAGs across documents of varying lengths (Abstract, Introduction, Full Paper) based on mean values of answer faithfulness, context recall, and context precision.
We also conducted a case study to evaluate the four RAG variants on a research paper with long text, but testing them on its abstract, introduction, and full text respectively. This approach ensures content consistency while varying document length (ranging from 255 to 16,475 tokens), allowing us to assess the scalability of CausalRAG.
For each of the three text materials, we constructed a separate graph, tested $n=20$ questions, and evaluated all four RAG variants across three key metrics. As mentioned earlier, we employed an LLM-based approach for graph construction, and the validity of these graphs—built from a business school marketing paper—was further verified by experts.
This case study also examines the retrieval differences among regular RAG, GraphRAG, and our proposed CausalRAG (as shown in Figure 4). In this example, a user uploads a long paper and asks: "How do different combinations of influence tactics impact the likelihood of winning a sales contract?" Expert validation confirms a clear answer—salespeople use influence tactics such as compliance-based and internalization-based strategies to attract buyers’ attention, which inherently increases their chances of securing a contract.
Analyzing the retrieval process, we observe that regular RAG retrieves semantically related content but fails to capture key information, leading to a vague and uninformative response. GraphRAG correctly retrieves relevant nodes like attention, but its community report process dilutes their importance by incorporating extraneous nodes, introducing bias despite factual grounding. In contrast, CausalRAG retrieves the initial relevant nodes and expands upon them by identifying causal paths essential for answering the query, ensuring a precise and causally grounded response.
While this example clarifies CausalRAG’s internal process, our full experimental results (Figure 5) demonstrate its scalability. CausalRAG consistently outperforms other RAG variants across different document lengths. Additionally, regular RAG performs well on shorter documents, even surpassing GraphRAG. However, as document length increases—such as with full papers—GraphRAG overtakes regular RAG, aligning with our analytical analysis.
CausalRAG exhibits robustness across varying context lengths by preserving access to discrete information while leveraging causal estimation for grounded retrieval. More importantly, an additional advantage—only partially reflected here—is that with a larger step size $s$, CausalRAG can trace long and complex causal relationships hidden in lengthy documents, generating well-founded responses—a capability that is challenging for existing RAG systems.
5.4 Parameter Study
Figure 6: Parameter study showing how different parameter choices (k and s) affect model performance.
We also tested the impact of different parameter combinations of $k$ and $s$ on CausalRAG (as shown in Figure 6). Using the average of our evaluation metrics, we observe a consistent trend: the performance of CausalRAG improves as $k$ and $s$ increase. Specifically, the performance rises from 0.534 at $k=s=1$ to 0.824 at $k=s=5$, aligning with intuitive expectations.
Notably, the improvement is more pronounced when increasing $k$ from 1 to 3, suggesting that retrieving additional context enhances reasoning quality. However, when $k\geq 4$, performance gains become less significant, indicating possible saturation due to information redundancy. Similarly, while increasing $s$ generally leads to better results, its effect diminishes at higher values of $k$, where retrieval is already extensive.
These results suggest an optimal trade-off between performance and computational efficiency. While the highest values ($k=5,s=5$) yield the best results, moderate settings such as $k=3,s=3$ still achieve competitive performance with lower retrieval costs. Future work could explore adaptive strategies to adjust these parameters dynamically based on query complexity.
5.5 Conclusion and Future work
We introduced CausalRAG, a framework that integrates causality into RAG to enhance performance. Our analytical and experimental analysis identified key limitations of regular RAG—loss of contextual integrity, reliance on semantic similarity over causal relevance, and low precision. By leveraging causal graphs, CausalRAG retrieves causally grounded context, outperforming baseline RAGs across diverse text lengths and domains. Additionally, it improves LLM explainability, enhances answer groundedness, and reduces hallucinations in generative responses. Future work includes benchmarking CausalRAG on domain-specific tasks.
Limitations
While CausalRAG enhances retrieval effectiveness by integrating causality, it has certain limitations. First, constructing causal graphs from unstructured text relies on LLM-based extraction, which may introduce extra costs, particularly in complex or ambiguous cases. Second, the computational cost of expanding and analyzing causal paths increases with larger documents, potentially impacting retrieval efficiency in extreme cases with a large number of tokens.